# The Matthew Effect of AI Programming Assistants: A Hidden Bias in Software Evolution

### v2

@@ Abstract
AI-assisted programming is rapidly reshaping software development, with large language models (LLMs) enabling new paradigms such as vibe coding and agentic coding. While prior works have focused on prompt design and code generation quality, the broader impact of LLM-driven development on the iterative dynamics of software engineering remains underexplored. In this paper, we conduct large-scale experiments on thousands of algorithmic programming tasks and hundreds of framework selection tasks to systematically investigate how AI-assisted programming interacts with the software ecosystem. Our analysis reveals a striking Matthew effect: the more popular a programming language or framework, the higher the success rate of LLM-generated code. The phenomenon suggests that AI systems may reinforce existing popularity hierarchies, accelerating convergence around dominant tools while hindering diversity and innovation. We provide a quantitative characterization of this effect and discuss its implications for the future evolution of programming ecosystems.
Keywords AI Programming Assistants $\cdot$ Large Language Models (LLMs) $\cdot$ Code Generation $\cdot$ Vibe Coding $\cdot$ Agentic Coding $\cdot$ Matthew Effect $\cdot$ Ecosystem Bias; Programming Language Diversity $\cdot$ Framework Popularity $\cdot$ Software Engineering Innovation $\cdot$ Benchmarking; LeetCode Evaluation $\cdot$ Framework-Level Analysis $\cdot$ Ecosystem Evolution $\cdot$ Bias Amplification.
@@ 1 Introduction
Large language models (LLMs) have quickly become ubiquitous in software engineering practice, with nearly all programmers (daigle2024survey, ) utilizing AI coding tools. Tools such as GitHub Copilot, Cursor, and integrated LLM-based coding assistants now support developers in algorithmic problem solving (yan2023closer, ), debugging, and even full-stack system construction. These advances introduce new coding paradigms: vibe coding, where developers iterate by prompting rather than typing every line, and agentic coding, where autonomous agents plan and execute end-to-end development tasks. Vibe coding democratizes software development (gadde2025democratizing, ) by lowering barriers to creation, translating conceptual intent into executable implementation. Agent-based code generation highlights the transformative potential of multi-agent systems in addressing the limitations of standalone LLMs. Agentic Coding effectively handles real-world coding challenges (wu2024autogen, ) by leveraging external tools for retrieval, achieving significant improvements (huang2023agentcoder, ) in pass rates across diverse benchmarks (zhang2024codeagent, ). Collectively, AI Coding could be the silver bullet for software engineering.
@@
Despite these breakthroughs, long-term ecosystem-level consequences of AI programming assistance remain underexplored. This research gap is critical because biases in training data and model behavior may systematically influence which languages, frameworks, and paradigms thrive or decline. Several observations underscore why this matters. First, LLM performance is uneven across languages: high-resource ecosystems such as Python achieve disproportionately strong results, while lower-resource languages receive much weaker support. For instance, the StarCoder dataset shows Python alone accounts for nearly 40% of its training corpus, while many other languages appear only marginally (li2023starcoder, ). Similarly, CodeGen explicitly notes that model quality varies substantially depending on training data availability, with mainstream ecosystems benefiting disproportionately (nijkamp2022codegen, ). Second, the bias extends beyond languages to frameworks and usage patterns. AI coding assistants often over-rely on established libraries, such as NumPy, which appears in up to 48% of completions even when alternatives may be more suitable, and they also display a persistent preference for Python, which is selected 58% of the time for performance-critical tasks where other languages like Rust may be objectively better (twist2025llms, ).Taken together, these patterns raise a central question: do AI tools genuinely empower innovation by lowering entry barriers, or do they inadvertently reinforce existing dominance hierarchies?
@@
The hypothesis we explore is that AI programming assistants exhibit a Matthew effect: “the rich get richer." This effect is rooted in the operational mechanics of LLMs, which are trained on massive datasets of publicly available code. Such dynamics risk creating lock-in effects that suppress experimentation and reduce opportunities for paradigm-shifting innovations, which is consistent with prior observations on programming language adoption and diffusion (meyerovich2013empirical, ). Programming learners may increasingly favor languages where AI support is strongest, further consolidating existing trends (prather2023robots, ). The empirical research on language adoption demonstrates that ecosystem factors (libraries, existing code, community size), rather than purely technical merit, strongly drive which languages gain traction, implying that model-mediated productivity gains could differentially amplify preexisting popularity patterns (meyerovich2013empirical, ). Most existing studies of AI-assisted code generation focus on short-term, micro-level evaluations that measure model performance on narrow benchmarks or single-language datasets, which do not capture the multi-faceted complexity of real-world software engineering. If this impact is overlooked, the resulting cycle, where popular languages receive more LLM support due to their prevalence in training data, risks reducing programming ecosystem diversity. Thus, AI assistance could simultaneously lower barriers to entry while stifling long-term innovation.
@@
  • We construct the first large-scale benchmark combining algorithmic programming tasks (Total $3011\times 8\times 5=120,440$) and complex full-stack development tasks to assess AI programming assistants across languages and frameworks.
@@
  • We provide the first empirical evidence that the Matthew effect emerges simultaneously at both the language and the framework levels in LLM code generation, showing how this dual-layer bias shapes software ecosystem trajectories and discussing its implications for innovation and diversity.
@@ 2.1 AI Programming Assistants
Research on LLM-based programming has largely focused on improving code generation quality, prompt engineering, and usability. Systems like Codex and Copilot have demonstrated high productivity gains in everyday development. More recent models such as GPT-4, Gemini, and DeepSeek exhibit strong reasoning and multi-step planning abilities, further lowering the barrier to complex programming. Recent work has examined the capabilities of AI-assisted code generation tools across diverse benchmarks. Early evaluations using HumanEval (chen2021evaluating, ) revealed that while Copilot often produced syntactically valid code, correctness rates were low and highly correlated with the prevalence of languages in the training data (yetistiren2022assessing, ). Although HumanEval became an early standard for evaluating LLM coding proficiency (jiang2024survey, ), its limited number of problems restricts its applicability across all research contexts. To address this limitation, subsequent studies employed LeetCode problems. Copilot’s best accuracy in Java (nguyen2022empirical, ), subsequent analysis extended to multiple tools, reporting that Copilot excelled in Java, ChatGPT maintained strong cross-language consistency (batista2024code, ), and Gemini performed best in JavaScript. Larger-scale evaluations showed Copilot’s accuracy decreased with problem difficulty and varied substantially across languages (mo2025assessing, ). The hyperscale multilingual benchmark-XCODEEVAL (khan2024xcodeeval, ), demonstrated persistent challenges for program synthesis and translation, especially in less common languages, though its automatically collected data lacks manual verification and may introduce noise. Existing benchmarks offer valuable insights into AI code generation, but most focus on a few popular languages. By contrast, we use a standardized LeetCode-based dataset to examine how language popularity affects performance.
@@ 2.2 Programming Ecosystem Evolution
The evolution of programming languages and frameworks has long been shaped by ecosystem factors such as community size, tooling, and industry adoption, often outweighing intrinsic technical merit (meyerovich2013empirical, ). In longitudinal study of web framework popularity over fifteen years, swacha2023evolution combined GitHub repository counts and Stack Overflow activity to demonstrate that adoption trajectories vary considerably across ecosystems. Recent surveys of LLMs for code highlight similar concerns: while these models show promise across a range of SE tasks, their performance is uneven and heavily biased toward widely used languages (zhang2023unifying, ). Human-curated benchmarks such as CodeArena reveal persistent cross-language disparities, suggesting that LLMs may reinforce the dominance of mainstream ecosystems like Python and JavaScript (yang2024evaluating, ). Beyond language adoption, ecosystem-level inequalities also emerge when LLMs act as autonomous coding agents: AgentBench (liu2023agentbench, ) shows that proprietary models such as GPT-4 exhibit superior reasoning and multi-turn decision-making, while open-source models lag significantly. Together, these findings point to a risk that AI programming assistants could entrench existing hierarchies and amplify long-term ecosystem imbalances.
@@
3 Environment and Benchmark Construction
3.1 Algorithmic tasks
3.1.1 Language Selection
We select eight languages guided (Table 1) by the June 2025 TIOBE Index (tiobe2025, ): Python, C++, Java, JavaScript, Go, Rust, Erlang, and Racket. These cover a spectrum from top-ranked mainstream languages to niche or emerging languages, enabling examination of popularity effects. This design allows us to investigate whether language popularity correlates with the performance of AI-assisted code generation and submission success.
@@ 3.1.2 Test Cases from LeetCode
To evaluate performance across these languages, we source algorithmic tasks from LeetCode (leetcode, ), an online platform chosen for its extensive collection of problems, robust online judging system, and broad multilingual support. LeetCode accommodates nearly 20 programming languages, including both widely used ones (C++, Java, Python) and less common ones (Rust, Scala, Elixir), making it an ideal environment for our multilingual study.
@@ 4.1 AI Coding
For each of the 3,011 problems crawled from LeetCode, we apply a standardized procedure wherein the problem statement and constraints are formatted into a consistent prompt template (for each of the eight selected programming languages). In total, this process results in over 120,440 individual code generation requests (3,011 problems × 8 languages × 5 models), by calling the APIs of these five models: GPT-4o-mini (hurst2024gpt, ), DeepSeek-V3 (liu2024deepseek, ), Gemini-2.0-Flash (Google_Gemini_2_Flash, ), Gemini-2.5-Flash (comanici2025gemini, ), Qwen3-Turbo (yang2025qwen3, ). This prompt is then submitted to each commercial closed-source LLM ’s API to generate solutions.
@@ 4.3 Results
Our large-scale evaluation reveals a pronounced performance gap between popular and less popular programming languages, a disparity that is substantial and consistent across all five state-of-the-art models tested. As shown in Table 3, mainstream languages including Python, JavaScript, Java, and C++ achieve Pass@1 rates exceeding 60% in top-performing models. In stark contrast, less popular languages such as Erlang and Racket struggle dramatically, with success rates often below 25% and sometimes approaching zero. For instance, the best-performing model (DeepSeek-V3) achieves 79.81% Pass@1 for Python but only 24.31% for Erlang and 20.82% for Racket. This pattern confirms that language popularity is a stronger predictor of AI coding success than model capability alone. This phenomenon, a pronounced Matthew effect in AI-assisted programming, becomes even more dramatic when stratified by problem difficulty. As illustrated in Figure 2, the performance gap widens substantially as complexity increases. For Easy problems, the difference between popular and niche languages ranges from 45 to 82 percentage points. This gap expands significantly to 58 to 95 points on Hard problems, indicating that the advantage of data-rich languages scales non-linearly with reasoning complexity. On these Hard tasks, top models achieve 50 to 63% success with popular languages but only 0 to 6% with less popular ones, demonstrating that superior model capability cannot compensate for the disadvantage of language unpopularity.
@@
Gemini-2.5-Flash
Python 67.92% 2045 217 0 726 23 609 1104 331
@@
Gemini-2.0-Flash
Python 62.94% 1895 268 0 787 61 594 1025 275
@@
Qwen3-Turbo
Python 37.00% 1114 401 0 1117 54 405 610 99
@@
Table 4: Statistical significance of Pass@1 differences between popular languages (Python, C++, Java, JavaScript) and less popular languages (Go, Rust, Erlang, Racket). All tests used paired t-tests across the 3,011 problems.
@@
DeepSeek-V3 +49.6 <0.001<0.001
Gemini-2.5-Flash +46.1 <0.001<0.001
Gemini-2.0-Flash +44.7 <0.001<0.001
GPT-4o-mini +38.0 <0.001<0.001
Qwen3-Turbo +34.2 <0.001<0.001
The observed Matthew effect has profound implications for programming language ecosystems. As AI-assisted programming becomes pervasive, the massive performance advantage for popular languages may accelerate their dominance while marginalizing niche languages, regardless of their technical merits. This could ultimately reduce linguistic diversity in software development. Our study uniquely leverages “data contamination" as a direct signal of language popularity, defining the overlap between test tasks and training data as a measure of a language’s representation in the training corpus. This premise dictated our deliberate selection of newly released LeetCode tasks from 2025. This choice minimizes “rote recall" from widely-circulated problems and aligns the “contamination gap" with contemporary popularity trends, allowing us to establish a clearer causal chain: language popularity dictates training data coverage, which in turn drives AI performance.
@@ 5.2 Results
Taken together, these findings illustrate a self-reinforcing feedback loop with significant implications for software ecosystem diversity: popular frameworks are easier for LLMs to generate successfully, developers relying on AI assistants are nudged toward these frameworks, and increased adoption further amplifies their online presence, ensuring even more model exposure in future iterations. Such dynamics exemplify the Matthew Effect in software ecosystems, where established technologies “get richer" in terms of visibility, usability, and adoption. While convenient for practitioners using mainstream stacks, this trend risks stifling ecosystem diversity by systematically disadvantaging technically promising but less popular frameworks. The findings further reveal that framework maturity and ecosystem support significantly impact AI code generation, with emerging frameworks lagging behind, suggesting that LLM-based assistance could exacerbate adoption gaps between established and new technologies.

### v3

@@ Abstract
AI-assisted programming is rapidly reshaping software development, with large language models (LLMs) enabling new paradigms such as vibe coding and agentic coding. While prior works have focused on prompt design and code generation quality, the broader impact of LLM-driven development on the iterative dynamics of software engineering remains underexplored. In this paper, we conduct large-scale experiments on thousands of algorithmic programming tasks and hundreds of framework selection tasks to systematically investigate how AI-assisted programming interacts with the software ecosystem. Our analysis quantifies a substantial performance asymmetry: mainstream languages and frameworks achieve significantly higher success rates than niche ones. This disparity suggests a feedback loop consistent with the Matthew Effect, where data-rich ecosystems gain superior AI support. While not the sole driver of adoption, current models introduce a non-negligible productivity friction for niche technologies, representing a hidden bias in software evolution.
@@ 1 Introduction
Large language models (LLMs) have quickly become ubiquitous in software engineering practice, with nearly all programmers (Daigle and Staff, 2024) utilizing AI coding tools. Tools such as GitHub Copilot, Cursor, and integrated LLM-based coding assistants now support developers in algorithmic problem solving (Yan et al., 2023), debugging, and even full-stack system construction. These advances introduce new coding paradigms: vibe coding, where developers iterate by prompting rather than typing every line, and agentic coding, where autonomous agents plan and execute end-to-end development tasks. Vibe coding democratizes software development (Gadde, 2025) by lowering barriers to creation, translating conceptual intent into executable implementation. Agent-based code generation highlights the transformative potential of multi-agent systems in addressing the limitations of standalone LLMs. Agentic Coding effectively handles real-world coding challenges (Wu et al., 2024) by leveraging external tools for retrieval, achieving significant improvements (Huang et al., 2023) in pass rates across diverse benchmarks (Zhang et al., 2024). Collectively, AI Coding could be the silver bullet for software engineering.
Prior to this empirical reality check, there was widespread optimism that LLMs would serve as a “Great Equalizer.” Recent studies show that LLMs help narrow the skill gap for junior developers (Noy and Zhang, 2023; Metabob, 2024). However, the belief that they also function as a language equalizer, making specific syntax irrelevant as suggested by (Huang, 2024) , has not yet been tested in empirical settings. We challenge this assumption and argue that instead of flattening the landscape, AI support introduces a critical new factor that may disproportionately disadvantage niche ecosystems.
@@
Long-term ecosystem-level consequences of AI programming assistance remain underexplored. This research gap is critical because biases in training data and model behavior may systematically influence which languages, frameworks, and paradigms thrive or decline. Several observations underscore why this matters. First, LLM performance is uneven across languages: high-resource ecosystems such as Python achieve disproportionately strong results, while lower-resource languages receive much weaker support. For instance, the StarCoder dataset shows Python alone accounts for nearly 40% of its training corpus, while many other languages appear only marginally (Li et al., 2023). Similarly, CodeGen explicitly notes that model quality varies substantially depending on training data availability, with mainstream ecosystems benefiting disproportionately (Nijkamp et al., 2022). Second, the bias extends beyond languages to frameworks and usage patterns. AI coding assistants often over-rely on established libraries, such as NumPy, which appears in up to 48% of completions even when alternatives may be more suitable, and they also display a persistent preference for Python, which is selected 58% of the time for performance-critical tasks where other languages like Rust may be objectively better (Twist et al., 2025).Taken together, these patterns raise a central question: do AI tools genuinely empower innovation by lowering entry barriers, or do they inadvertently reinforce existing dominance hierarchies?
@@
The hypothesis we explore is that AI programming assistants exhibit a Matthew effect: “the rich get richer.” This effect is rooted in the operational mechanics of LLMs, which are trained on massive datasets of publicly available code. Such dynamics risk creating lock-in effects that suppress experimentation and reduce opportunities for paradigm-shifting innovations, which is consistent with prior observations on programming language adoption and diffusion (Meyerovich and Rabkin, 2013). Programming learners may increasingly favor languages where AI support is strongest, further consolidating existing trends (Prather et al., 2023). The empirical research on language adoption demonstrates that ecosystem factors (libraries, existing code, community size), rather than purely technical merit, strongly drive which languages gain traction, implying that model-mediated productivity gains could differentially amplify preexisting popularity patterns (Meyerovich and Rabkin, 2013). Most existing studies of AI-assisted code generation focus on short-term, micro-level evaluations that measure model performance on narrow benchmarks or single-language datasets, which do not capture the multi-faceted complexity of real-world software engineering. If this impact is overlooked, the resulting cycle, where popular languages receive more LLM support due to their prevalence in training data, risks reducing programming ecosystem diversity. Thus, AI assistance could simultaneously lower barriers to entry while stifling long-term innovation.
@@
  • We construct the first large-scale benchmark combining algorithmic programming tasks (Total $3011\times 9\times 5=135,495$) and complex full-stack development tasks to assess AI programming assistants across languages and frameworks.
@@
  • We quantify a substantial asymmetry in AI support across languages and frameworks, demonstrating patterns consistent with Matthew-effect dynamics. While separating AI-specific amplification from pre-existing structural biases remains an open empirical question, our findings reveal a measurable “AI productivity tax” that correlates with ecosystem popularity.
@@ 2.1 AI Programming Assistants
Research on LLM-based programming has largely focused on improving code generation quality, prompt engineering, and usability. Systems like Codex and Copilot have demonstrated high productivity gains in everyday development. More recent models such as GPT-4, Gemini, and DeepSeek exhibit strong reasoning and multi-step planning abilities, further lowering the barrier to complex programming. Recent work has examined the capabilities of AI-assisted code generation tools across diverse benchmarks. Early evaluations using HumanEval (Chen et al., 2021) revealed that while Copilot often produced syntactically valid code, correctness rates were low and highly correlated with the prevalence of languages in the training data (Yetistiren et al., 2022). Although HumanEval became an early standard for evaluating LLM coding proficiency (Jiang et al., 2024), its limited number of problems restricts its applicability across all research contexts. To address this limitation, subsequent studies employed LeetCode problems. Copilot’s best accuracy in Java (Nguyen and Nadi, 2022), subsequent analysis extended to multiple tools, reporting that Copilot excelled in Java, ChatGPT maintained strong cross-language consistency (Batista et al., 2024), and Gemini performed best in JavaScript. Larger-scale evaluations showed Copilot’s accuracy decreased with problem difficulty and varied substantially across languages (Mo et al., 2025). The hyperscale multilingual benchmark-XCODEEVAL (Khan et al., 2024), demonstrated persistent challenges for program synthesis and translation, especially in less common languages, though its automatically collected data lacks manual verification and may introduce noise. Existing benchmarks offer valuable insights into AI code generation, but most focus on a few popular languages. By contrast, we use a standardized LeetCode-based dataset to examine how language popularity affects performance.
@@ 2.2 Programming Ecosystem Evolution
The evolution of programming languages and frameworks has long been shaped by ecosystem factors such as community size, tooling, and industry adoption, often outweighing intrinsic technical merit (Meyerovich and Rabkin, 2013). In a longitudinal study of web framework popularity, Swacha and Kulpa (2023) demonstrated that adoption trajectories vary considerably across ecosystems. AI as a New Adoption Factor: The emergence of AI assistants introduces a complex new variable to this evolutionary dynamic. Recent literature presents a nuanced picture of AI-assisted development: while studies report substantial speedups on specific tasks (Peng et al., 2023), concerns regarding correctness, over-reliance, and workflow disruption persist (Weisz et al., 2024). This suggests that AI is not a simple linear accelerator but a nuanced influence on developer experience. However, recent surveys highlight a critical risk within this influence: LLM performance is uneven and heavily biased toward widely used languages (Zhang et al., 2023). Human-curated benchmarks such as CodeArena reveal persistent cross-language disparities, suggesting that LLMs may reinforce the dominance of mainstream ecosystems like Python and JavaScript (Yang et al., 2024). Beyond language adoption, ecosystem-level inequalities also emerge when LLMs act as autonomous coding agents: AgentBench (Liu et al., 2023) shows that proprietary models such as GPT-4 exhibit superior reasoning and multi-turn decision-making, while open-source models lag significantly. Together, these findings point to a risk that AI programming assistants could entrench existing hierarchies and amplify long-term ecosystem imbalances.
3 Environment and Benchmark Construction
3.1 Algorithmic tasks
3.1.1 Language Selection
We select nine languages guided (Table 1) by the June 2025 TIOBE Index (TIOBE Software BV, 2025): Python, C++, C, Java, JavaScript, Go, Rust, Erlang, and Racket. These cover a spectrum from top-ranked mainstream languages to niche or emerging languages, enabling examination of popularity effects. This design allows us to investigate whether language popularity correlates with the performance of AI-assisted code generation and submission success.
@@ 3.1.2 Test Cases from LeetCode
To evaluate performance across these languages, we source algorithmic tasks from LeetCode (LeetCode, 2025), an online platform chosen for its extensive collection of problems, robust online judging system, and broad multilingual support. LeetCode accommodates nearly 20 programming languages, including both widely used ones (C++, Java, Python) and less common ones (Rust, Scala, Elixir), making it an ideal environment for our multilingual study.
@@
We position these algorithmic tasks as the ‘canary in the coal mine’ for linguistic competency. While acknowledging that LeetCode does not capture the full software ecosystem, the high compile error rates observed in niche languages expose a fundamental deficiency. If an LLM cannot generate syntactically correct code for basic logic, this failure inherently precludes its effective application in broader and more complex engineering contexts.
@@ 4.1 AI Coding
For each of the 3,011 problems crawled from LeetCode, we apply a standardized procedure wherein the problem statement and constraints are formatted into a consistent prompt template (for each of the nine selected programming languages). In total, this process results in over 135,495 individual code generation requests (3,011 problems × 9 languages × 5 models), by calling the APIs of these five models: GPT-4o-mini (Hurst et al., 2024), DeepSeek-V3 (Liu et al., 2024), Gemini-2.0-Flash (Google, 2025), Gemini-2.5-Flash (Comanici et al., 2025), Qwen3-Turbo (Yang et al., 2025). This prompt is then submitted to each commercial closed-source LLM ’s API to generate solutions.
@@ 4.3 Results
Our large-scale evaluation reveals a pronounced performance gap between popular and less popular programming languages, a disparity that is substantial and consistent across all five state-of-the-art models tested. As shown in Table 3, mainstream languages including Python, JavaScript, Java, C and C++ achieve Pass@1 rates exceeding 60% in top-performing models. In stark contrast, less popular languages such as Erlang and Racket struggle dramatically, with success rates often below 25% and sometimes approaching zero. For instance, the best-performing model (DeepSeek-V3) achieves 79.81% Pass@1 for Python but only 24.31% for Erlang and 20.82% for Racket. This pattern confirms that language popularity is a stronger predictor of AI coding success than model capability alone. This phenomenon, a pronounced Matthew effect in AI-assisted programming, becomes even more dramatic when stratified by problem difficulty. As illustrated in Figure 2, the performance gap widens substantially as complexity increases. For Easy problems, the difference between popular and niche languages ranges from 45 to 82 percentage points. This gap expands significantly to 58 to 95 points on Hard problems, indicating that the advantage of data-rich languages scales non-linearly with reasoning complexity. On these Hard tasks, top models achieve 50 to 63% success with popular languages but only 0 to 6% with less popular ones, demonstrating that superior model capability cannot compensate for the disadvantage of language unpopularity.
@@
Gemini-2.5-Flash Python 67.92% 2045 217 0 726 23 609 1104 331
@@
Gemini-2.0-Flash Python 62.94% 1895 268 0 787 61 594 1025 275
@@
Qwen3-Turbo Python 37.00% 1114 401 0 1117 54 405 610 99
@@
Table 4: Statistical significance of Pass@1 differences between popular languages (Python, C++, C, Java, JavaScript) and less popular languages (Go, Rust, Erlang, Racket). All tests used paired t-tests across the 3,011 problems.
@@
DeepSeek-V3 +44.8 <0.001<0.001
Gemini-2.5-Flash +42.3 <0.001<0.001
Gemini-2.0-Flash +40.5 <0.001<0.001
GPT-4o-mini +33.1 <0.001<0.001
Qwen3-Turbo +28.9 <0.001<0.001
The observed Matthew effect has profound implications for programming language ecosystems. As AI-assisted programming becomes pervasive, the massive performance advantage for popular languages may accelerate their dominance while marginalizing niche languages, regardless of their technical merits. This could ultimately reduce linguistic diversity in software development. Our study uniquely leverages “data contamination” as a direct signal of language popularity, defining the overlap between test tasks and training data as a measure of a language’s representation in the training corpus. This premise dictated our deliberate selection of newly released LeetCode tasks from 2025. This choice minimizes “rote recall” from widely-circulated problems and aligns the “contamination gap” with contemporary popularity trends, allowing us to establish a clearer causal chain: language popularity dictates training data coverage, which in turn drives AI performance.
@@ 5.2 Results
Taken together, these findings illustrate a self-reinforcing feedback loop with significant implications for software ecosystem diversity: popular frameworks are easier for LLMs to generate successfully, developers relying on AI assistants are nudged toward these frameworks, and increased adoption further amplifies their online presence, ensuring even more model exposure in future iterations. Such dynamics exemplify the Matthew Effect in software ecosystems, where established technologies “get richer” in terms of visibility, usability, and adoption. While convenient for practitioners using mainstream stacks, this trend risks stifling ecosystem diversity by systematically disadvantaging technically promising but less popular frameworks. The findings further reveal that framework maturity and ecosystem support significantly impact AI code generation, with emerging frameworks lagging behind, suggesting that LLM-based assistance could exacerbate adoption gaps between established and new technologies.